The following information relates to Windsor Inc.’s accounts receivable for the 2020 fiscal year: 1. An aging schedule of the accounts receivable as at December 31, 2020, is as follows: Age Net Debit Balance % to Be Applied after Write-off Is Made Under 60 days $173,000 2% 61–90 days 142,000 4% 91–120 days 44,000 * 8% Over 120 days 22,000 $5,200 definitely uncollectible; 20% of remainder is estimated uncollectible $381,000 *The $2,800 write off of receivables (see item 4 below) is related to the 91–120-day category. 2. The Accounts Receivable control account has a debit balance of $381,000 on December 31, 2020. 3. Two entries were made in the Bad Debt Expense account during the year: (1) a debit on December 31 for the amount credited to Allowance for Doubtful Accounts, and (2) a credit for $2,800 on November 3, 2020, and a debit to Allowance for Doubtful Accounts because of a bankruptcy. 4. Allowance for Doubtful Accounts is as follows for 2020: Allowance for Doubtful Accounts 11/3 Uncollectible accounts written off 2,800 1/1 Beginning balance 8,750 12/31 6% of $381,000 22,860 5. There is a credit balance in Accounts Receivable (61–90 days) of $4,900, which represents an advance on a sales contract. Assuming that the books have not been closed for 2020, make the necessary correcting entries. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.) Account Titles and Explanation Debit Credit (To correct entry for account written off) (To correct entry for advance on sales contract) (To record uncollectible account) (To adjust allowance to year-end computed balance)

Account Titles and Explanation

Debit

Credit

(To correct entry for account written off)

(To correct entry for advance on sales contract)

(To record uncollectible account)

(To adjust allowance to year-end computed balance)
Answer #1

Requirement: Prepare the following journal entries

Date Account Title and Explanation Debit Credit
Dec. 31 Bad Debt Expense $2,800
                 Accounts Receivable $2,800
To correct entry for the accounts receivable write off
Dec. 31 Accounts Receivable $4,900
                 Unearned Sales Revenue $4,900
To correct entry for the entry on sale contract
Dec. 31 Allowance for Doubtful Accounts $5,200
                Accounts Receivable $5,200
To record uncollectible accounts written off
Dec. 31 Allowance for Doubtful Accounts $7,618
                Bad Debt Expense $7,618
To adjust the allowance for doubtful accounts

Notes: Compute the adjustment required as follows

Particulars Amount
Unadjusted balance in allowance for doubtful accounts $8,750
Add: 6% of $381,000 $22,860
Deduct: Uncollectibles written off ($2,800 + $5,200) ($8,000)
Balance available $23,610
Deduct: Allowance for doubtful accounts required ($15,992)
Excess allowance for doubtful amounts $7,618

Compute required allowance on doubtful accounts as follows

Particulars Amount × Rate = Amount
Under 60 days $173,000 × 2% = $3,460
61- 90 ($142,000 + $4,900) $146,900 × 4% = $5,876
91-120 ($44,000 − $2,800) $41,200 × 8% = $3,296
Over 120 days ($22,000 − $5,200) $16,800 × 20% = $3,360
Total $377,900 $15,992